OpenStudy (anonymous):

Find the coordinates of the other endpoint of a segment with the given endpoint and mid point 28) Q(-5, 9) M(-8, -2) 29) A(6,7) M(10, -7) 30) D(-3.5, -6) M(1.5, 4.5)

OpenStudy (anonymous):

28.P(-11,-13) 29.B(14,-21) 30.C(6.5,15)

OpenStudy (anonymous):

Can you show the work for these problems? Please

OpenStudy (anonymous):

use mid point formula if end points are (x,y)and (a,b) let (p,q) be the mid point then p=(a+x/2,b+y/2)

OpenStudy (anonymous):

How would it look like?

OpenStudy (anonymous):

take second question let (x,y) be an end point then from given data 10=x+6/2 -7=y+7/2

OpenStudy (anonymous):

OHHHHHHHHH!!!! Okay so it would be -5=x+9/2 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and then the bottom one is going to be -8=y+(-2)/2?

OpenStudy (anonymous):

How did you get -11 for the first point? I am getting -19